OpenStudy (anonymous):

differentiate y= sin^-1(x^2)

OpenStudy (anonymous):

lol i just figured it out :)

OpenStudy (anonymous):

and, i hope this helps \[\frac{d}{dx}[\arcsin u] = \frac{u \prime}{\sqrt{1-u^2}}\]

OpenStudy (anonymous):

oh alright lol

OpenStudy (anonymous):

thanks :)